A cyclist riding a bike with a gear ratio of 4 pedals with a speed of 55 rpm . If the radius of the bicycle's wheels is 31cm, what is the translational speed of the bicycle as it moves?
Answer #1

here,

angular speed of pedaling , w1 = 55 rpm

angular speed of rear wheel , w2 = 55 * 4

w2 = 220 rpm

w2 = 220 * 2pi/60 = 23.03 rad/s

radius , r = 0.31 m

translational speed of the bicycle = r * w2

translational speed of the bicycle = 0.31 * 23.03

translational speed of the bicycle = 7.14 m/s

the translational speed of the bicycle is 7.14 m/s